# The Symmetry Taco: Equivalences between Gapped, Gapless, and Mixed-State SPTs

### v1

@@ IV.3 Example: $\mathbb{Z}_{4}\times\mathbb{Z}_{4}$ Folded SPT
The first two terms impose the energetic constraints Eq. (32) on the two copies of the system and correspond to the condensation of $e^{2}m^{2}$ and $\bar{e}^{2}\bar{m}^{2}$ on each copy in the corresponding $\mathcal{D}(\mathbb{Z}_{4})$ and $\mkern 1.5mu\overline{\mkern-1.5mu\mathcal{D}(\mathbb{Z}_{4})\mkern-1.5mu}\mkern 1.5mu$ SymTFTs. The second line constructs the SymTaco  from the decoupled layers and corresponds to tunneling $e^{2}$ between the two copies (equivalently, condensing $e^{2}\bar{e}^{2}$), and spontaneously breaks the original $\mathbb{Z}_{4}\times\mathbb{Z}_{4}$ symmetry to $\mathbb{Z}_{2}^{A}\times\mathbb{Z}_{4}^{AB}$. The final term fully gaps out the system and corresponds to tunneling $em$ and $e^{3}m$ between the two copies. The ground of this Hamiltonian realizes the non-trivial $1+1$d gapped $\mathbb{Z}_{4}\times\mathbb{Z}_{2}$ SPT that was discussed as Example 3 in Sec. II.2.
V Mixed-State SymTFT: Folded Boundaries as Mixed-State SPTs
Figure 4: SymTFT for the Choi state. (a) Distinction between strong and weak symmetry operators. (b) Distinction between strong and weak order parameters. The bold red lines connecting the layers represent decoherence ($a_{+}\bar{a}_{-}$ anyon condensation), which has the effect of coupling the ket and bra spaces of |$\ket{\rho}\rangle$.
@@ V Mixed-State SymTFT: Folded Boundaries as Mixed-State SPTs
A natural question then is whether the SymTFT picture can be extended to mixed states, accommodating their aforementioned novel symmetry structure and describing intrinsically mixed-state phases. In this Section, we answer this in the affirmative: given a $1+1$d $G$-symmetric mixed state $\rho$, we propose the SymTaco  as the SymTFT for the corresponding Choi state, a vectorized version of $\rho$. Using this framework, we arrive at an exact correspondence between iASPTs, igSPTs, and the folded SPTs discussed in previous Sections. We emphasize that, while this correspondence between different classes of SPTs builds on each of the prior Sections, the only background required for readers solely interested in the SymTFT description for mixed-state SPTs is provided in Section II.
@@ V.2 Choi Isomorphism and Mixed-State Correlations
In the analysis of mixed states, it is convenient to express the operators $O$ acting on $\mathcal{H}$ as a vectors |$\ket{O}\rangle$ in a “doubled” Hilbert space using the Choi-Jamiołkowski isomorphism [^12] We will follow the standard convention in the mixed-state literature and refer to this vectorization of $\rho$ as the “Choi state” |$\ket{\rho}\rangle$. Strictly speaking, the Choi state refers to the vectorized version of a quantum channel, not a density matrix. [104, 105]. This allows us to map the problem of characterizing density matrices to that of characterizing pure states (subject to the constraints of Hermiticity and positive semi-definiteness). For instance, the density matrix $\rho$ is mapped to a state |$\ket{\rho}\rangle\in\mathcal{H}_{+}\otimes\mathcal{H}_{-}$ as
@@
$U_{k+}\ket{\rho}\rangle=U_{k-}^{*}\ket{\rho}\rangle=e^{i\theta}\ket{\rho}\rangle\,.$ (54)
Hence, |$\ket{\rho}\rangle$ describes a bilayer state that is symmetric under the diagonal action of $G\subset G_{+}\times G-$ as well as the action of $K\subset G$ on either layer. The action of the symmetries, in the SymTFT picture described in the following Section, is depicted in Fig. 4.
@@
In the following subsection, we propose a SymTFT for describing “gapped” mixed-state phases. For pure states, the properties of the ground state of a local Hamiltonian are strongly constrained by the latter’s spectral gap. In $1+1$d, gapped, non-degenerate ground states of local Hamiltonians exhibit exponentially decaying correlations–with the correlation length set by the gap–and are short-range entangled [110]. The existence of a gap allows for well-defined notions of phases of matter, such that gapped ground states within the same phase are related by quasi-adiabatic Hamiltonian deformations. In principle, one can also define phases of matter based solely on states rather than Hamiltonians, with the appropriate equivalence relation provided by quasi-local finite-depth unitary evolution (locality, as encoded in a Lieb-Robinson bound, is the crucial ingredient for distinguishing phases based on their long-range correlations). In the mixed-state setting, there is no canonical notion of a “gap” and hence no canonical notion of a phase. Despite this, there is considerable ongoing effort in making precise the notion of a mixed-state phase, primarily by exploring different equivalence relations on states [111, 112, 113] and using (emergent) anomalies to characterize mixed-state phases [71, 72, 75, 73].
In what follows, we will adopt a perspective analogous to that employed in e.g., Refs. [74, 75], and use the Choi state |$\ket{\rho}\rangle$, to define the properties of the mixed-state $\rho$. In particular, we will say that $\rho$ is short-range entangled if |$\ket{\rho}\rangle$ is short-range entangled; likewise, we will define $\rho$ as a “gapped” mixed state if |$\ket{\rho}\rangle$ is a gapped pure state (i.e., has exponentially decaying correlations of local observables or is the ground state of a–fictitious–gapped Hamiltonian in the doubled Hilbert space). While this simplification appears dramatic, it has nonetheless proven extremely useful in diagnosing mixed-state phases as it permits one to leverage known results regarding pure state phases to characterize mixed states. An important caveat, though, is that under this simplification, we do not expect the characterization of mixed-state phases to coincide with other proposals (see Ref. [74] for a discussion regarding this point) [^14] For instance, the $\mathbb{Z}_{2}$ SWSSB state discussed in Sec. V.4 is manifestly separable, but would be considered long-range entangled as per this definition, since the corresponding Choi state is a GHZ state in the doubled space.. That being said, defining mixed-state phase via their Choi states constitutes a perfectly valid, consistent definition of mixed-state phases and is one that is naturally suited in the context of the SymTFT correspondence we develop in the following subsection, where we propose the Choi states of pure-state TOs as the SymTFTs for symmetric mixed states.
@@ V.3 Mixed-State SymTFT and the Symmetry Taco
In this subsection, we propose a SymTFT framework for one-dimensional gapped phases of mixed states with total symmetry group $G$. Recall that we take a mixed state $\rho$ to be gapped if its Choi state |$\ket{\rho}\rangle$ describes a gapped pure state. We will argue that phases of mixed states with $G_{+}\times G_{-}$ symmetry are classified by the data governing allowed anyon condensations in a bulk topological order $\mathcal{D}(G)\boxtimes\overline{\mathcal{D}(G)}$ that satisfy a symmetry and positivity constraint. These constraints implement the Hermiticity and positivity conditions that mixed states are required to satisfy. We find that these constraints are identical to the constraints characterizing folded SPTs (see Sec. IV), which is central to the tetraptych of correspondences shown in Fig. 1 and establishes the SymTaco as the appropriate bulk object for studying $G$-symmetric mixed states. Indeed, folding the $\mathcal{D}(G)$ TO along $\mathcal{D}(G)/\mathcal{A}$ can physically be interpreted as maximal decoherence of the anyons in $\mathcal{D}(G)/\mathcal{A}$, which corresponds a diagonal condensate of anyons between the ket and bra layers of $\mathcal{D}(G)/\mathcal{A}\boxtimes\overline{\mathcal{D}(G)/\mathcal{A}}$, as shown in Fig. 5. This perspective reproduces the classification of mixed-state SPTs in one dimension and provides a transparent method for extracting the relevant string order parameters.
@@
We begin by considering the case where the $1+1$d mixed state has a strong $G$ symmetry. The bulk TO encoding the symmetry data of the boundary ASPT is given by the bilayer TO $\mathcal{D}(G)\boxtimes\overline{\mathcal{D}(G)}$. This is to be understood as a topological order in the doubled space, arising from the density matrix of a pure state $\mathcal{D}(G)$ TO. We will label anyons of the bilayer TO as $a_{+}b_{-}$, where $a,b\in\mathcal{D}(G)$; here, the second anyon is always understood to have opposite (complex conjugate) spin and braiding. Similarly to the pure state case discussed in Sec. II, we realize the $1+1$d mixed-state SPT as a gapped boundary to vacuum of the $\mathcal{D}(G)\boxtimes\overline{\mathcal{D}(G)}$ using the thin slab construction. As before, the thin slab has two boundaries, one of which we take to be the symmetry boundary and the other the dynamical boundary. On the symmetry boundary, we condense all pure charges of the TO. This corresponds to the Lagrangian algebra $\text{Rep}(G_{+}\times G_{-})$. The line operators for the confined anyons form the category $\text{Vec}(G_{+}\times G_{-})$, which label the symmetry operators acting on the mixed state [^15] As in the previous subsection, we have used $G_{+}$ and $G_{-}$ to label the “ket” and “bra” symmetries of the mixed state.. The strong symmetries are given by the confined fluxes that live entirely within a single layer, while the weak symmetries are given by diagonal fluxes which have support on both layers, as we will see in later examples (see Fig. 4(a)).
@@
Just as in the pure state case, the $1+1$d state in the Choi space is specified by a pattern of anyon condensation on the dynamical boundary. However, the allowed anyon condensates are now constrained by the fact that the Choi state must describe a density matrix, which is Hermitian and positive semi-definite. We claim that these constraints restrict the allowed anyon condensates precisely to those describing the folded boundaries discussed in Section IV.1. We will show this explicitly in the Abelian case and hypothesize that this correspondence holds as well in the non-Abelian case.
Let us begin with the Abelian case. Suppose that $\rho_{0}=\ket{\psi_{0}}\bra{\psi_{0}}$ | describes a $2+1$d pure state hosting Abelian topological order, $\mathcal{C}$. Its corresponding Choi state |$\ket{\rho_{0}}\rangle$ then describes a bilayer topological order, $\mathcal{C}\times\overline{\mathcal{C}}$. Suppose now that we subject $\rho_{0}$ to a local finite-depth quantum channel $\mathcal{E}[\cdot]$, which can be understood as forcing the incoherent proliferation of anyon excitations. The resulting density matrix $\rho\equiv\mathcal{E}[\rho_{0}]$ then describes a mixed-state topological order, as discussed in Refs. [71, 72]. The Choi state |$\ket{\rho}\rangle$ of this decohered density matrix again describes a topological order in the doubled Hilbert space.
In the doubled space, the effect of the decoherence on the original state |$\ket{\rho_{0}}\rangle$ is to induce the conventional condensation of bulk anyons [114]. In the following, for ease of comparison with the discussion of folded boundaries, we use the notation $(a,b)\sim a_{+}b_{-}$. Let us define the indicator function $M_{(a,b)}$ as in Eq. (34) to keep track of which anyons from the original TO are condensed, where $a$ ($b$) lives in the ket (bra) space: $(a,b)\in\mathcal{C}\times\overline{\mathcal{C}}$. Now, for an anyon $a$, we can write down a Wilson line operator $W_{a}^{\gamma}$ on an open string $\gamma$, which creates $a$ and its conjugate $\bar{a}$ at its endpoints. An anyon $a$ is condensed if $W_{a}^{\gamma}$ exhibits long-range correlation as a function of the distance between the end-points of the open string $\gamma$. For ease of presentation, we restrict ourselves to “fixed-point” states in which the anyons have been maximally decohered [71]–in this case an anyon $a$ is condensed if the application of the open string operator $W_{a}^{\gamma}$ leaves $\rho$ unchanged. Suppose $(a,\bar{b})$ is condensed; then,
@@
by Hermiticity, so $(b,\bar{a})$ is also condensed. Thus, if $M_{(a,b)}=1$, then we must have $M_{(\bar{b},\bar{a})}=M_{(b,a)}=1$, where the second equality follows since the condensed anyons form a group. Hence, we see that Hermiticity leads to the condition that anyon condensation is symmetric between the two layers.
@@
where we used cyclicity of the trace and the definition of the inner product on the doubled space. Now, using the Cauchy-Schwarz inequality, we find that, if $(a,\bar{b})$ is condensed,
@@
The two traces in the final line must both be non-zero in order to satisfy the inequality, implying that both $(a,\bar{a})$ and $(b,\bar{b})$ must also be condensed. Using the fact that the condensed anyons form a group, this in turn implies that $(a\times\bar{b},1)$ must also be condensed. Hence, if $M_{(a,\bar{b})}=1$, we have that $M_{(a\times\bar{b},1)}=1$ as well.
@@
Finally, let us turn to the case where $\rho$ has a strong symmetry $K\lhd G$ and the total symmetry group $G$ is an extension of the weak symmetry by $K$, as in Eq. (51) (we do not restrict to Abelian $K$ or $G$ here). This is to contrast with the previous situation where $\rho$ has a $G_{+}\times G_{-}$ symmetry, which could be spontaneously broken down to such a subgroup after decoherence. Here, we wish to realize the symmetry breaking explicitly, via condensation of electric anyons in the bulk Choi state TO. As noted in Sec. V.1, the full symmetry group in the Choi space is $K\rtimes G$, consisting of elements of the form $(kg,g)$. In this case, the bulk TO in the Choi space will be isomorphic to $\mathcal{D}(K\rtimes G)$. To obtain this TO, we can begin with the original $\mathcal{D}(G_{+}\times G_{-})$ TO and and condense charges to break the gauge group down to $K\rtimes G$. Let $\bigoplus_{i}\pi_{i}\subset\mathcal{D}(G)$ be the set of charges condensed such that the resulting TO is given by $\mathcal{D}(K)$, for $K\subset G$. These charges correspond to those that are uncharged under $K$. Then, condensing the charges $\bigoplus_{i}\pi_{i}\pi^{*}_{i}\subset\mathcal{D}(G_{+}\times G_{-})$ leads to a bulk topological order $\mathcal{D}(K\rtimes G)$ in the Choi space. This condensation is positive, since the anyons condensed are of the form $a_{+}\bar{a}_{-}$. Physically, this condensation corresponds to starting from a pure state $\mathcal{D}(G)$ TO and decohering the $\pi_{i}$ charges, which can be done by applying maximal decoherence composed of the short string operators for these charges (see Refs. [71, 72] for the explicit procedure). Hence, the bulk TO of a $1+1$d theory with a $K\subset G$ strong symmetry is given by a mixed state TO whose Choi state topological order is isomorphic to $\mathcal{D}(K\rtimes G)$.
@@ V.4 Example: Strong-to-weak SSB
|$\ket{\mathrm{GHZ}}$ $=\frac{1}{\sqrt{2}}\left(\ket{0}^{\otimes L}+\ket{1}^{\otimes L}\right)\,,$ (66)
@@
Let us see how this physics is reproduced in the SymTaco. The bulk TO in the Choi space for a strong $\mathbb{Z}_{2}$ symmetry is givne by $\mathcal{D}(\mathbb{Z}_{2}^{+}\times\mathbb{Z}_{2}^{-})$, which is the Choi state of the pure state $\mathbb{Z}_{2}$ Toric Code. The anyons in the Choi space form a $\mathbb{Z}_{2}^{4}$ group under fusion, generated by the charges $e_{+},e_{-}$ and the fluxes $m_{+},m_{-}$, with a $-1$ braiding between $e_{+}$ and $m_{+}$, as well as between $e_{-}$ and $m_{-}$. The reference boundary of the SymTaco is chosen to be the charge condensed boundary, with the Lagrangian subgroup given by $\mathcal{L}=\langle e_{+},e_{-}\rangle$. With this choice, the $m_{+}$ and $m_{-}$ generate the $\mathbb{Z}_{2}^{+}\times\mathbb{Z}_{2}^{-}$ symmetry, respectively. The weak symmetry, corresponding to the diagonal subgroup $\mathbb{Z}_{2}\subset\mathbb{Z}_{2}^{+}\times\mathbb{Z}_{2}^{-}$ is then generated by the composite flux $m_{+}m_{-}$. Due to the nontrivial braiding of the charges with fluxes, we see that the local order parameters for breaking of the full symmetry $\mathbb{Z}_{2}^{+}\times\mathbb{Z}_{2}^{-}$ correspond to the Wilson lines of $e_{+}$ and $e_{-}$ stretching from the reference boundary to the dynamical boundary. The order parameter for breaking down to the diagonal subgroup $\mathbb{Z}_{2}$ is then given by the Wilson line for $e_{+}e_{-}$ which, indeed, has trivial braiding with the generator of this preserved symmetry, $m_{+}m_{-}$. In the original physical space, this corresponds to the order parameter for SWSSB.

### v2

@@ V Mixed-State SymTFT: Folded Boundaries as Mixed-State SPTs
A natural question then is whether the SymTFT picture can be extended to mixed states, accommodating their aforementioned novel symmetry structure and describing intrinsically mixed-state phases. In this Section, we answer this in the affirmative: given a 1+11+1d $G$-symmetric mixed state $\rho$, we propose the SymTaco  as the SymTFT for the corresponding Choi state, a vectorized version of $\rho$. Using this framework, we arrive at an exact correspondence between iASPTs, igSPTs, and the folded SPTs discussed in previous Sections. We emphasize that, while this correspondence between different classes of SPTs builds on each of the prior Sections, the only background required for readers solely interested in the SymTFT description for mixed-state SPTs is provided in Section II.
@@ V.2 Choi Isomorphism and Mixed-State Correlations
[…]
@@
[…]
@@
In the following subsection, we propose a SymTFT for describing “gapped” mixed-state phases. For pure states, the properties of the ground state of a local Hamiltonian are strongly constrained by the latter’s spectral gap. In 1+11+1d, gapped, non-degenerate ground states of local Hamiltonians exhibit exponentially decaying correlations–with the correlation length set by the gap–and are short-range entangled [114]. The existence of a gap allows for well-defined notions of phases of matter, such that gapped ground states within the same phase are related by quasi-adiabatic Hamiltonian deformations. In principle, one can also define phases of matter based solely on states rather than Hamiltonians, with the appropriate equivalence relation provided by quasi-local finite-depth unitary evolution (locality, as encoded in a Lieb-Robinson bound, is the crucial ingredient for distinguishing phases based on their long-range correlations). In the mixed-state setting, there is no canonical notion of a “gap” and hence no canonical notion of a phase. Despite this, there is considerable ongoing effort in making precise the notion of a mixed-state phase, primarily by exploring different equivalence relations on states [115, 116, 117] and using (emergent) anomalies to characterize mixed-state phases [74, 75, 80, 76].
[…]
@@ V.3 Mixed-State SymTFT and the Symmetry Taco
[…]
@@
We begin by considering the case where the 1+11+1d mixed state has a strong $G$ symmetry. The bulk TO encoding the symmetry data of the boundary ASPT is given by the bilayer TO $\mathcal{D}(G)\boxtimes\overline{\mathcal{D}(G)}$. This is to be understood as a topological order in the doubled space, arising from the density matrix of a pure state $\mathcal{D}(G)$ TO. We will label anyons of the bilayer TO as $a_{+}b_{-}$, where $a,b\in\mathcal{D}(G)$; here, the second anyon is always understood to have opposite (complex conjugate) spin and braiding. Similarly to the pure state case discussed in Sec. II, we realize the 1+11+1d mixed-state SPT as a gapped boundary to vacuum of the $\mathcal{D}(G)\boxtimes\overline{\mathcal{D}(G)}$ using the thin slab construction. As before, the thin slab has two boundaries, one of which we take to be the symmetry boundary and the other the dynamical boundary. On the symmetry boundary, we condense all pure charges of the TO. This corresponds to the Lagrangian algebra $\text{Rep}(G_{+}\times G_{-})$. The line operators for the confined anyons form the category $\text{Vec}(G_{+}\times G_{-})$, which label the symmetry operators acting on the mixed state [^15] As in the previous subsection, we have used $G_{+}$ and $G_{-}$ to label the “ket” and “bra” symmetries of the mixed state.. The strong symmetries are given by the confined fluxes that live entirely within a single layer, while the weak symmetries are given by diagonal fluxes which have support on both layers, as we will see in later examples (see Fig. 4(a)).
@@
Just as in the pure state case, the 1+11+1d state in the Choi space is specified by a pattern of anyon condensation on the dynamical boundary. However, the allowed anyon condensates are now constrained by the fact that the Choi state must describe a density matrix, which is Hermitian and positive semi-definite. We claim that these constraints restrict the allowed anyon condensates precisely to those describing the folded boundaries discussed in Section IV.1. We will show this explicitly in the Abelian case and hypothesize that this correspondence holds as well in the non-Abelian case.
[…]
@@
by Hermiticity, so $b_{+}\bar{a}_{-}$ is also condensed. Thus, if $M_{a,b}=1$, then we must have $M_{\bar{b},\bar{a}}=M_{b,a}=1$, where the second equality follows since the condensed anyons form a group. Hence, we see that Hermiticity leads to the condition that anyon condensation is symmetric between the two layers.
@@
where we used cyclicity of the trace and the definition of the inner product on the doubled space. Now, using the Cauchy-Schwarz inequality, we find that, if $a_{+}\bar{b}_{-}$ is condensed,
@@
The two traces in the final line must both be non-zero in order to satisfy the inequality, implying that both $a_{+}\bar{a}_{-}$ and $b_{+}\bar{b}_{-}$ must also be condensed. Using the fact that the condensed anyons form a group, this in turn implies that $(a\times\bar{b})_{+}1_{-}$ must also be condensed. Hence, if $M_{a,\bar{b}}=1$, we have that $M_{a\times\bar{b},1}=1$ as well.
@@
Finally, let us turn to the case where $\rho$ has a strong symmetry $K\lhd G$ and the total symmetry group $G$ is an extension of the weak symmetry by $K$, as in Eq. (51) (we do not restrict to Abelian $K$ or $G$ here). This is to contrast with the previous situation where $\rho$ has a $G_{+}\times G_{-}$ symmetry, which could be spontaneously broken down to such a subgroup after decoherence. Here, we wish to realize the symmetry breaking explicitly, via condensation of electric anyons in the bulk Choi state TO. As noted in Sec. V.1, the full symmetry group in the Choi space is $K\rtimes G$, consisting of elements of the form $(kg,g)$. In this case, the bulk TO in the Choi space will be isomorphic to $\mathcal{D}(K\rtimes G)$. To obtain this TO, we can begin with the original $\mathcal{D}(G_{+}\times G_{-})$ TO and and condense charges to break the gauge group down to $K\rtimes G$. Let $\bigoplus_{i}\pi_{i}\subset\mathcal{D}(G)$ be the set of charges condensed such that the resulting TO is given by $\mathcal{D}(K)$, for $K\subset G$. These charges correspond to those that are uncharged under $K$. Then, condensing the charges $\bigoplus_{i}\pi_{i}\pi^{*}_{i}\subset\mathcal{D}(G_{+}\times G_{-})$ leads to a bulk topological order $\mathcal{D}(K\rtimes G)$ in the Choi space. This condensation is positive, since the anyons condensed are of the form $a_{+}\bar{a}_{-}$. Physically, this condensation corresponds to starting from a pure state $\mathcal{D}(G)$ TO and decohering the $\pi_{i}$ charges, which can be done by applying maximal decoherence composed of the short string operators for these charges (see Refs. [74, 75] for the explicit procedure). Hence, the bulk TO of a 1+11+1d theory with a $K\subset G$ strong symmetry is given by a mixed state TO whose Choi state topological order is isomorphic to $\mathcal{D}(K\rtimes G)$.
@@ V.4 Example: Strong-to-weak SSB
[…]
@@
Let us see how this physics is reproduced in the SymTaco. The bulk TO in the Choi space for a strong $\mathbb{Z}_{2}$ symmetry is givne by $\mathcal{D}(\mathbb{Z}_{2}^{+}\times\mathbb{Z}_{2}^{-})$, which is the Choi state of the pure state $\mathbb{Z}_{2}$ Toric Code. The anyons in the Choi space form a $\mathbb{Z}_{2}^{4}$ group under fusion, generated by the charges $e_{+},e_{-}$ and the fluxes $m_{+},m_{-}$, with a 1-1 braiding between $e_{+}$ and $m_{+}$, as well as between $e_{-}$ and $m_{-}$. The reference boundary of the SymTaco is chosen to be the charge condensed boundary, with the Lagrangian subgroup given by $\mathcal{L}=\langle e_{+},e_{-}\rangle$. With this choice, the $m_{+}$ and $m_{-}$ generate the $\mathbb{Z}_{2}^{+}\times\mathbb{Z}_{2}^{-}$ symmetry, respectively. The weak symmetry, corresponding to the diagonal subgroup $\mathbb{Z}_{2}\subset\mathbb{Z}_{2}^{+}\times\mathbb{Z}_{2}^{-}$ is then generated by the composite flux $m_{+}m_{-}$. Due to the nontrivial braiding of the charges with fluxes, we see that the local order parameters for breaking of the full symmetry $\mathbb{Z}_{2}^{+}\times\mathbb{Z}_{2}^{-}$ correspond to the Wilson lines of $e_{+}$ and $e_{-}$ stretching from the reference boundary to the dynamical boundary. The order parameter for breaking down to the diagonal subgroup $\mathbb{Z}_{2}$ is then given by the Wilson line for $e_{+}e_{-}$ which, indeed, has trivial braiding with the generator of this preserved symmetry, $m_{+}m_{-}$. In the original physical space, this corresponds to the order parameter for SWSSB.